# Attention Pruning: Automated Fairness Repair of Language Models via Surrogate Simulated Annealing

### v2

@@ 8. Conclusion
In this paper, we tackle the problem of mitigating unfairness in pre-trained LLMs by carefully pruning attention heads. Although attention head pruning is a computationally hard problem, we found that techniques based on surrogate simulated annealing are effective and scalable to improve the fairness of LLMs while preserving the utility. We implemented our approach, Attention Pruning, and showed that it outperforms the state-of-the-art post-processing bias mitigator by identifying an ideal subset of heads to prune.
For future work, developing genetic algorithms that explores the top $k$ states found during the search, rather than selecting the best state, is a promising direction. Another direction is to extend the head pruning idea to other responsible AI requirements, such as privacy and robustness.

### v3

@@ 8. Conclusion
In this paper, we tackle the problem of mitigating unfairness in pre-trained LLMs by carefully pruning attention heads. Although attention head pruning is a computationally hard problem, we found that techniques based on surrogate simulated annealing are an effective and scalable strategy to improve the fairness of LLMs while preserving the utility.
We implemented our approach, Attention Pruning (AP), and showed that it outperforms the state-of-the-art post-processing bias mitigator by identifying an ideal subset of heads to prune. AP achieves up to 40% reduction in gender bias and outperforms the state-of-the-art approaches by identifying attention head configurations with both low bias and low perplexity.
For future work, developing genetic algorithms that explore the top $k$ states found during the search, rather than selecting the best state, is a promising direction. Another direction is to extend the head pruning idea to other responsible AI requirements, such as privacy and robustness.
Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. CNS-2527657 and CNS-2230061.